sunisround:

Jessica invites some friends to a cowboy game. she purchases a suite at the stadium for $1,500. She also orders a $250 pizza hut package. Three other family members already had tickets to the game, but join for the pizza. If the prices were divided evenly between Jessica and her friends - and the cost was $57.58 per person - how many friends did Jessica invite to the game?

sunisround:

I want to know how to set up the problem.

sunisround:

like what equation do i use?

Vocaloid:

Hm. My intuition would be that you would add the cost of the suite + pizza package then divide by x and set that equal to the price per person

Vocaloid:

(1500+250)/x = 57.58

Vocaloid:

but that doesn't give me a whole number so that can't be it

sunisround:

what about the "three" family members?

Vocaloid:

the wording is throwing me off

sunisround:

i did not make it up, my teacher gave it to me and he probably modified it with different words so we would not cheat.

Vocaloid:

yeah it's not your fault, let me try and work something out

sunisround:

okay thank you ill wait patiently.

Vocaloid:

this is my best attempt the 1500 suite is split between Jessica and her friends the 250 pizza is split between Jessica, her friends, and her family so if we divide total cost/total people for each part, we get 57.78

Vocaloid:

let x = the number of friends

Vocaloid:

1500/(Jessica + friends) + 250/(Jessica+friends+ 3 family members) = 57.78 1500(x+1) + 250/(x+1+3) = 57.78 this gives me x ~ 30

sunisround:

so jessica is x right?

Vocaloid:

x is the number of friends

sunisround:

Jessica + Friends= x+1

Vocaloid:

yeah, friends is x and Jessica is 1 person

sunisround:

ok i get it know because the problem wants us to find the number of friends right

Vocaloid:

yes

sunisround:

but wait why did you multiply 1500(x+1) and not divide?

Vocaloid:

there should be a division sign between them sorry

sunisround:

ok

sunisround:

i want to do this by hand and not by calculator how will i start with (1500/x+1)+(250/x+1+3)=57.78?

sunisround:

does 250/x+1+3 simplify to 250/x+4 right?

Vocaloid:

yes

Vocaloid:

you would just make of the denominators the same by finding the lcd

Vocaloid:

*both of the denominators

sunisround:

how do i do that?

Vocaloid:

your denominator on the left is (x+1) the denominator on the right is (x+4) you would multiply the left denominator by (x+4) and the right denominator by (x+1)

Vocaloid:

that will make both denominators (x+1)(x+4)

Vocaloid:

then, you have to perform the same operation to the numerators to keep the fractions equivalent

sunisround:

so is it 1500+250/(x+1)(x+4)?

Vocaloid:

you also have to perform the same operations to the numerators

sunisround:

so (x+1)(x+4)* the numerators?

Vocaloid:

1500(x+4)


(x+4)(x+1) 250(x+1)
(x+1)(x+4)

Vocaloid:

pretend those lines are fraction bars

sunisround:

okay.

Vocaloid:

then add them together

sunisround:

wait i am confused.

sunisround:

why didnt you multiply 1500 time x+1?

sunisround:

times*

Vocaloid:

1500/(x+1) + 250/(x+4) are your original fractions you multiply the numerator and the denominator of 1500/(x+1) by (x+4) to convert the denominator to (x+1)(x+4) there would be no point in multiplying it by (x+1) since that's already in the denominator

sunisround:

ok

sunisround:

so like...

sunisround:

(x+4) 1500 (x+1) 250



+
---- =57.78 (x+4) (x+1) (x+1) (x+4)

Vocaloid:

@sillybilly123 can you tell me if I am interpreting this problem correctly?

sunisround:

wow this thing came out ugly lol.

Vocaloid:

can you try using the "equation" button perhaps?

sunisround:

ok sorry i am new

sunisround:

let me try and solve it

sillybilly123:

missing info unless you make assumptions

Vocaloid:

yeah, I made my best attempt based on my interpretation of the problem

sunisround:

\[(1750x+6250)/(x+1)(x+4)\]

sillybilly123:

i mean, Jessica is a "person" and she cannot end up paying a flat rate as she paid for the suite

sunisround:

i know but i guess my teacher does not like to have good grammar or something is wrong with his printer.

sunisround:

so what do i do next?

Vocaloid:

you would set that equal to 57.58 and solve for x

sunisround:

multiply (x+1)(x+4) times 57.78? and then solve for x?

Vocaloid:

(1750x+6250) = 57.58(x+1)(x+4) if that's what you mean

sunisround:

yes, is that correct?

Vocaloid:

yes

Vocaloid:

expand the Right hand side using foil and then combine like terms

sunisround:

yeah

Vocaloid:

I would leave out the 57.58 until the very end after you have finished expanding (x+1)(x+4)

Vocaloid:

then you would multiply every term of the expansion by 57.58

sunisround:

ok

sunisround:

1745x+6250=x^2?

sunisround:

6246*

sunisround:

1745x+6246=x^2

Vocaloid:

I'm not exactly sure what you're doing ;-;

sunisround:

you said to multiply 57.78 at the end

Vocaloid:

start by expanding this (x+1)(x+4)

sunisround:

1750x+6250=x^2+5x+4?

Vocaloid:

1750x+6250=(57.58)(x^2+5x+4)

sunisround:

ok

Vocaloid:

then distribute the 57.58

sunisround:

1750x+6250=57.78x^2+288.9x+231.12?

Vocaloid:

good now combine like terms

sunisround:

they are already combined, or are they not?

sunisround:

i just need to solve for x right?

Vocaloid:

move everything to one side of the equation

sunisround:

ok

Vocaloid:

then apply the quadratic formula

sunisround:

-7869.667713?

sunisround:

to have that amount of friends she must be on drugs or something

Vocaloid:

let's back up a bit

sunisround:

ok to where?

Vocaloid:

to here: 1750x+6250=57.78x^2+288.9x+231.12

Vocaloid:

subtract 1750x from both sides

sunisround:

ok wait let me erase something because my paper is getting full

sunisround:

0=57.78^2-1461.1x-6018.88?

Vocaloid:

yeah now apply the quadratic formula

sunisround:

okay

sunisround:

x~28.893?

sunisround:

so 29?

Vocaloid:

yup

Vocaloid:

[it's a bit off 29 because of rounding error]

Vocaloid:

anyway, I have no guarantees this is the answer but again this was my best attempt

sunisround:

alright thank you very much anyways, i probably gave you a headache, because i am having one. have a great day.